Grade 12th passPhysical Chemistry

If 250cm3 of 0.2M kcl and250cm3 of 0.2M k2so4 are combined then what would be the molar concentration of K+ in the resulting solution. Also mention the method of solving such questions related to solutions.

August
10 Years agoGrade 12th pass

1 Answer

Umakant biswal
10 Years ago
@ august 
first calculate the volume . 
if u will calculate the volume u will get it as 500 cc .. 
after that find the total no.of K+ ions.
then calculate by the following formula 
volume*molarity*n factor 
and acc to the question by calculating 
u will get it as 
K+ = 250*0.2*1 + 250*0.2*2 = 150
the molarity will be 0.2 in each case here .. 
then for finding the final answer apply the following formula 
conc of K+ = no.of K+/volume
by applying the above formula u will get 
 150/500 = 0.3M
which will be the answer 
HOPE U WILL GET HOW TO SOLVE SUCH QUESTIONS 
ALL THE BEST 
IF MORE DOUBT THEN ASK .